# Measuring Faithfulness Depends on How You Measure: Classifier Sensitivity in LLM Chain-of-Thought Evaluation

### 20 Mar 2026 v1

@@ Abstract
Recent work on chain-of-thought (CoT) faithfulness reports single aggregate numbers (e.g., DeepSeek-R1 acknowledges hints 39% of the time [1]), implying that faithfulness is an objective, measurable property of a model. This paper demonstrates that it is not. Three classifiers (a regex-only detector, a two-stage regex-plus-LLM pipeline, and an independent Claude Sonnet 4 judge) are applied to 10,276 influenced reasoning traces from 12 open-weight models spanning 9 families and 7B to 1T parameters. On identical data, these classifiers produce overall faithfulness rates of 74.4%, 82.6%, and 69.7%, respectively. Per-model gaps range from 2.6 to 30.6 percentage points. At the hint-type level, all pairwise McNemar tests are significant ($p<0.001$). The disagreements are systematic, not random: inter-classifier agreement measured by Cohen’s $\kappa$ ranges from 0.06 (“slight”) for sycophancy hints to 0.42 (“moderate”) for grader hints, and the asymmetry is pronounced: for sycophancy, 883 cases are classified as faithful by the pipeline but unfaithful by the Sonnet judge, while only 2 go the other direction. Classifier choice can also reverse model rankings: Qwen3.5-27B ranks 1st under the pipeline but 7th under the Sonnet judge; OLMo-3.1-32B moves in the opposite direction, from 9th to 3rd. The root cause is that different classifiers operationalize related faithfulness constructs at different levels of stringency (lexical mention versus epistemic dependence), and these constructs yield divergent measurements on the same behavior. These results demonstrate that published faithfulness numbers cannot be meaningfully compared across studies that use different classifiers, and that future evaluations should report sensitivity ranges across multiple classification methodologies rather than single point estimates.
@@ 1 Introduction
This paper quantifies the problem. Three classifiers (a regex-only detector, a two-stage regex-plus-LLM pipeline, and an independent Claude Sonnet 4 judge) are applied to 10,276 identical cases where 12 open-weight reasoning models [15] changed their answers in response to injected hints. Three results challenge the practice of reporting single faithfulness numbers:
@@ 3 Experimental Setup
The dataset is drawn from a larger study of CoT faithfulness in open-weight reasoning models. It consists of 498 questions (300 from MMLU [25] and 198 from GPQA Diamond [26]), each paired with five hint types: sycophancy, consistency, metadata, grader, and unethical, the same categories used by Chen et al. [1]. Each hint steers the model toward a specific incorrect answer. Twelve open-weight reasoning models spanning 7B to 1T parameters across 9 model families are tested, including DeepSeek-R1 [15], OLMo [18], and Qwen-family models [17]. From the full set of hinted runs, the 10,276 influenced cases, instances where the model changed its answer from the baseline to match the hinted answer, are extracted, and all three classifiers are applied to this identical set.
@@
    Regex-only. The first stage of the pipeline in isolation: 38 regular expression patterns spanning all five hint types, designed for high precision. A match means the CoT explicitly references the hint (e.g., “the professor suggested,” “according to the metadata”). Free and instantaneous, but limited to surface-level lexical matching.
@@
    Regex + Ollama judges (full pipeline). Stage 1 regex as above, followed by Stage 2: a three-judge LLM panel (GLM-5, Kimi K2, Gemini 3 Flash) running locally via Ollama. Cases not caught by regex receive majority-vote classification. Effectively zero marginal cost.
@@
    Claude Sonnet 4 (independent judge). Applied to all 10,276 cases via the Anthropic Batch API, following the LLM-as-judge paradigm [12]. Intended to assess whether the hint served as a reason for the model’s answer, not merely mentioned in passing. Cost: \$48.99 for the full dataset. Binary YES/NO output.
@@ 4.2 Disagreements Are Hint-Type Specific
Figure 2 visualizes the per-hint-type rates side by side. The sycophancy gap (+43.4 percentage points) is 15 times larger than the grader gap (+2.9 percentage points). Classifier choice barely matters for grader and unethical hints but dominates the measurement for sycophancy and consistency. This pattern aligns with Sharma et al.’s [16] finding that sycophantic behaviors are particularly difficult to operationalize, as they exist on a continuum between social acknowledgment and epistemic dependence. Hu et al. [29] similarly found that sycophancy in reasoning models requires real-time calibration rather than binary classification. Metadata is the only hint type where Sonnet is more generous than the pipeline ($\Delta=-9.5$ pp), suggesting that the Sonnet judge identifies implicit metadata references that the regex patterns miss.
@@ 4.3 Formal Agreement Statistics
Raw percentage agreement between the pipeline and Sonnet classifiers can be misleading when base rates are imbalanced (e.g., when one classifier labels nearly all cases as faithful). Cohen’s $\kappa$ corrects for chance agreement and provides a more interpretable measure of inter-classifier reliability. Table 4 reports $\kappa$ by hint type. Agreement ranges from “slight” ($\kappa=0.06$) for sycophancy to “moderate” ($\kappa=0.42$) for grader, reinforcing the finding that classifier disagreement concentrates on specific hint types rather than distributing uniformly. The low $\kappa$ for sycophancy and consistency reflects the “mention vs. dependence” ambiguity discussed below: both classifiers detect the same surface cues but assign them different epistemic weight, a pattern consistent with Jacovi and Goldberg’s [7] argument that faithfulness is construct-dependent and with the causal ambiguity highlighted by Lanham et al. [4].
Table 4: Cohen’s $\kappa$ and McNemar’s $\chi^{2}$ between the pipeline and Sonnet classifiers, by hint type. Interpretation bands follow Landis and Koch (1977): <<0.20 = slight, 0.21–0.40 = fair, 0.41–0.60 = moderate. All McNemar tests are significant at $p<0.001$.
@@ 4.4 Model Rankings Can Reverse
This has direct consequences for the field. A paper using a regex-based classifier would conclude that Qwen3.5-27B is the most faithful open-weight reasoner (rank 1); a paper using a Sonnet judge would rank it 7th of 12. Similarly, OLMo-3.1-32B ranks 3rd by Sonnet but 9th by the pipeline. Both sets of conclusions would be presented as properties of the models, when in fact they are partly properties of the measurement instrument, a concern that parallels Turpin et al.’s [3] observation that measured unfaithfulness depends on the evaluation methodology.
@@
Both readings are defensible. The regex approach operationalizes faithfulness as mention: did the CoT reference the hint at all? The Sonnet judge operationalizes it as dependence: did the hint serve as a reason for the answer, as opposed to being acknowledged and then overridden by independent reasoning? As Jacovi and Goldberg [7] argued, faithfulness admits multiple valid definitions; the present data show that these definitional choices have large empirical consequences. Zaman and Srivastava [30] reached a complementary conclusion, demonstrating that CoT can be faithful without explicitly verbalizing the hint, precisely the gap between the “mention” and “dependence” constructs observed here. This mirrors the broader challenge in CoT evaluation identified by Chua et al. [23], who noted that different operationalizations of faithfulness can yield contradictory conclusions.
This distinction explains the hint-type pattern in Table 2. Sycophancy hints (“a professor says…”) frequently appear in CoT traces that also contain independent verification, creating maximum ambiguity between mention and dependence. Grader hints (“the grading system expects…”) tend to be either fully embraced or fully ignored, leaving little room for interpretive disagreement.
@@ 5 Discussion
The central finding of this study is that CoT faithfulness, as currently operationalized in the literature, is not a single measurable quantity but a family of related constructs whose values depend critically on the classification methodology employed. This conclusion was anticipated on theoretical grounds by Jacovi and Goldberg [7], who argued that faithfulness definitions are inherently construct-dependent, and by Bean et al. [11], who identified analogous construct validity failures across 445 LLM benchmarks. The present results provide direct empirical confirmation: three classifiers intended to operationalize related faithfulness constructs at different levels of stringency, applied to the same 10,276 cases, produce faithfulness rates spanning a 12.9-percentage-point range overall and a 43.4-percentage-point range for sycophancy hints alone.
The divergence arises because the classifiers implement different operationalizations of the shared criterion. A regex classifier operationalizes faithfulness as mention: the CoT contains a lexical reference to the hint. An LLM judge [12] operationalizes it as epistemic dependence: the hint served as a causally load-bearing reason for the model’s conclusion, rather than being acknowledged and then overridden by independent analysis. These are distinct constructs that happen to share the same label. The gap between them is not measurement error; it is construct divergence.
This pattern has precedent outside faithfulness evaluation. Gordon et al. [9] demonstrated that correcting for annotator disagreement in toxicity classification reduced apparent model performance from 0.95 to 0.73 ROC AUC, a degradation comparable in magnitude to the gaps observed here. Plank [10] argued more broadly that such label variation reflects genuine interpretive plurality rather than noise, and should be modeled rather than suppressed. The present study extends these insights from human annotation to automated classification of reasoning traces, showing that the same interpretive ambiguity that produces inter-annotator disagreement in subjective tasks also produces inter-classifier disagreement in faithfulness evaluation.
@@
These results motivate approaches that are less sensitive to classification methodology. Baker et al. [37] argued that CoT may be highly informative despite being unfaithful at the surface level, Yang et al. [38] investigated the monitorability of CoT in reasoning models, and Xiong et al. [39] measured faithfulness specifically in thinking drafts. In a companion paper, the gap between a model’s internal thinking tokens and its visible CoT is found to provide a more robust signal of unfaithful reasoning, as thinking-token analysis does not require subjective judgments about whether a hint reference is “load-bearing.” Classifier-independent signals may ultimately prove more reliable than any single faithfulness metric [40].

### 23 Mar 2026 v2

@@ Abstract
Recent work on chain-of-thought (CoT) faithfulness reports single aggregate numbers (e.g., DeepSeek-R1 acknowledges hints 39% of the time [4]), implying that faithfulness is an objective, measurable property of a model. This paper provides evidence that it is not. Faithfulness is defined operationally as whether a model’s chain-of-thought acknowledges the injected hint as a factor in its answer. Three classifiers (a regex-only detector, a two-stage regex-plus-LLM pipeline, and an independent Claude Sonnet 4 judge) are applied to 10,276 influenced reasoning traces from 12 open-weight models spanning 9 families and 7B to 1T parameters. On identical data, these classifiers produce overall faithfulness rates of 74.4%, 82.6%, and 69.7%, respectively. Per-model gaps range from 2.6 to 30.6 percentage points. At the hint-type level, all pairwise McNemar tests are significant ($p<0.001$). The disagreements are systematic, not random: inter-classifier agreement measured by Cohen’s $\kappa$ ranges from 0.06 (“slight”) for sycophancy hints to 0.42 (“moderate”) for grader hints, and the asymmetry is pronounced: for sycophancy, 883 cases are classified as faithful by the pipeline but unfaithful by the Sonnet judge, while only 2 go the other direction. Classifier choice can also reverse model rankings: Qwen3.5-27B ranks 1st under the pipeline but 7th under the Sonnet judge; OLMo-3.1-32B moves in the opposite direction, from 9th to 3rd. The root cause is that different classifiers operationalize related faithfulness constructs at different levels of stringency (lexical mention versus epistemic dependence), and these constructs yield divergent measurements on the same behavior. These results indicate that published faithfulness numbers cannot be meaningfully compared across studies that use different classifiers, and that future evaluations should report sensitivity ranges across multiple classification methodologies rather than single point estimates.
@@ 1 Introduction
This paper provides an initial quantification of the problem. Three classifiers (a regex-only detector, a two-stage regex-plus-LLM pipeline, and an independent Claude Sonnet 4 judge) are applied to 10,276 identical cases where 12 open-weight reasoning models [7] changed their answers in response to injected hints. While the underlying data are drawn from a companion empirical study [39], the present paper addresses a distinct research question: not how faithful are these models, but how stable are faithfulness measurements across classification methodologies. This question requires a different analytical framework (inter-classifier agreement, construct comparison) than the cross-family analysis of the parent study. Four results challenge the practice of reporting single faithfulness numbers:
@@
Three hypotheses are tested:
H1:
Aggregate faithfulness rates differ significantly across the three classifiers when applied to the same data, reflecting construct-level differences rather than random measurement error.
H2:
The magnitude of classifier disagreement varies by hint type, with social-pressure hints (sycophancy, consistency) producing larger gaps than rule-based hints (grader, unethical).
H3:
Classifier choice can reverse the ordinal ranking of models, such that conclusions about which models are “most faithful” depend on the measurement instrument.
@@ 3 Experimental Setup
The dataset is drawn from a larger study of CoT faithfulness in open-weight reasoning models [39]. It consists of 498 questions (300 from MMLU [12] and 198 from GPQA Diamond [26]), each paired with five hint types: sycophancy, consistency, metadata, grader, and unethical, the same categories used by Chen et al. [4]. Each hint steers the model toward a specific incorrect answer. Twelve open-weight reasoning models spanning 7B to 1T parameters across 9 model families are tested, including DeepSeek-R1 [7], OLMo [21], and Qwen-family models [24]. All inference runs were conducted between February 25 and March 10, 2026, via the OpenRouter API using temperature=0.0=0.0 and seed=103=103 for reproducibility. Exact OpenRouter model identifiers, generation parameters, and regex patterns are documented in the companion study [39]. From the full set of hinted runs, the 10,276 influenced cases, instances where the model changed its answer from the baseline to match the hinted answer, are extracted, and all three classifiers are applied to this identical set.
@@
    Regex-only. The first stage of the pipeline in isolation: 38 regular expression patterns (listed in full in the companion study’s appendix [39]) spanning all five hint types, designed for high precision. A match means the CoT explicitly references the hint (e.g., “the professor suggested,” “according to the metadata”). Free and instantaneous, but limited to surface-level lexical matching.
@@
    Regex + Ollama judges (full pipeline). Stage 1 regex as above, followed by Stage 2: a three-judge LLM panel (GLM-5, Kimi K2, Gemini 3 Flash) running locally via Ollama. Cases not caught by regex receive majority-vote classification. All judge models use temperature=0.0=0.0 and max_tokens=2000=2000; thinking text is truncated to 4,000 characters before submission. Effectively zero marginal cost.
@@
    Claude Sonnet 4 (independent judge). Applied to all 10,276 cases via the Anthropic Batch API, following the LLM-as-judge paradigm [43]. Intended to assess whether the hint served as a reason for the model’s answer, not merely mentioned in passing (model version claude-sonnet-4-20250514, accessed via OpenRouter). Cost: \$48.99 for the full dataset. Binary YES/NO output.
@@ 4 Results
This section reports findings for each hypothesis in turn, followed by supplementary analyses of construct divergence and rank stability.
@@ 4.2 Disagreements Are Hint-Type Specific
Figure 2 visualizes the per-hint-type rates side by side. The sycophancy gap (+43.4 percentage points) is 15 times larger than the grader gap (+2.9 percentage points). Classifier choice barely matters for grader and unethical hints but dominates the measurement for sycophancy and consistency. Metadata is the only hint type where Sonnet is more generous than the pipeline ($\Delta=-9.5$ $p$).
@@ 4.3 Formal Agreement Statistics
Raw percentage agreement between the pipeline and Sonnet classifiers can be misleading when base rates are imbalanced (e.g., when one classifier labels nearly all cases as faithful). Cohen’s $\kappa$ corrects for chance agreement and provides a more interpretable measure of inter-classifier reliability. Table 4 reports $\kappa$ by hint type. Agreement ranges from “slight” ($\kappa=0.06$) for sycophancy to “moderate” ($\kappa=0.42$) for grader, confirming that classifier disagreement concentrates on specific hint types rather than distributing uniformly.
Table 4: Cohen’s $\kappa$ and McNemar’s $\chi^{2}$ between the pipeline and Sonnet classifiers, by hint type. Interpretation bands follow Landis and Koch [15]: <<0.20 = slight, 0.21–0.40 = fair, 0.41–0.60 = moderate. All McNemar tests are significant at $p<0.001$.
@@ 4.4 Model Rankings Can Reverse
Under the pipeline, Qwen3.5-27B would be ranked the most faithful open-weight reasoner (rank 1); under the Sonnet judge, it would rank 7th of 12. Similarly, OLMo-3.1-32B ranks 3rd by Sonnet but 9th by the pipeline.
@@
Both readings are defensible. The regex approach operationalizes faithfulness as mention: did the CoT reference the hint at all? The Sonnet judge operationalizes it as dependence: did the hint serve as a reason for the answer, as opposed to being acknowledged and then overridden by independent reasoning? The examples above illustrate the two constructs; their theoretical and practical implications are addressed in the Discussion.
Summary of findings.
All three hypotheses are supported. Overall faithfulness rates span a 12.9-percentage-point range across classifiers (H1). Per-hint-type gaps range from 2.9pp (grader) to 43.4pp (sycophancy), confirming that social-pressure hints produce the largest disagreements (H2). Model rankings reverse between classifiers, with Spearman $\rho=0.67$ indicating moderate but imperfect rank agreement (H3).
@@ 5 Discussion
This study set out to determine whether faithfulness measurements are stable across classification methodologies. The central finding is that they are not: three classifiers applied to the same 10,276 influenced reasoning traces produce overall faithfulness rates spanning a 12.9-percentage-point range, per-hint-type gaps as large as 43.4 percentage points, and model rankings that reverse depending on the classifier used. CoT faithfulness, as currently operationalized in the literature, does not appear to be a single measurable quantity but rather a family of related constructs whose values depend critically on the measurement instrument. This conclusion was anticipated on theoretical grounds by Jacovi and Goldberg [14], who argued that faithfulness definitions are inherently construct-dependent, and by Bean et al. [3], who identified analogous construct validity failures across 445 LLM benchmarks. The present results provide direct empirical confirmation at scale.
The divergence arises because the classifiers implement different operationalizations of the shared criterion. A regex classifier operationalizes faithfulness as mention: the CoT contains a lexical reference to the hint. An LLM judge [43] operationalizes it as epistemic dependence: the hint served as a causally load-bearing reason for the model’s conclusion, rather than being acknowledged and then overridden by independent analysis. These are distinct constructs that happen to share the same label. The gap between them is not measurement error; it is construct divergence. The hint-type pattern reinforces this interpretation: sycophancy hints exist on a continuum between social acknowledgment and epistemic dependence [28], creating maximum ambiguity between the two constructs and producing the largest classifier gaps. Hu et al. [13] similarly found that sycophancy in reasoning models resists binary classification and requires real-time calibration. Grader hints, by contrast, tend to be either fully embraced or fully ignored, leaving little room for interpretive disagreement. As Jacovi and Goldberg [14] argued, faithfulness admits multiple valid definitions; the present data show that these definitional choices have large empirical consequences. Zaman and Srivastava [41] reached a complementary conclusion, demonstrating that CoT can be faithful without explicitly verbalizing the hint—precisely the gap between the “mention” and “dependence” constructs observed here.
This pattern has precedent outside faithfulness evaluation. Gordon et al. [9] demonstrated that correcting for annotator disagreement in toxicity classification reduced apparent model performance from 0.95 to 0.73 ROC AUC, a degradation comparable in magnitude to the gaps observed here. Plank [23] argued more broadly that such label variation reflects genuine interpretive plurality rather than noise, and should be modeled rather than suppressed. The present study extends these insights from human annotation to automated classification of reasoning traces, suggesting that a similar interpretive ambiguity that produces inter-annotator disagreement in subjective tasks also produces inter-classifier disagreement in faithfulness evaluation.
@@
These results motivate approaches that are less sensitive to classification methodology. Baker et al. [32] argued that CoT may be highly informative despite being unfaithful at the surface level, Yang et al. [35] investigated the monitorability of CoT in reasoning models, and Xiong et al. [34] measured faithfulness specifically in thinking drafts. A companion study of thinking-answer divergence [40] finds that the gap between a model’s internal thinking tokens and its visible CoT may provide a more robust signal of unfaithful reasoning, because thinking-token analysis does not require subjective judgments about whether a hint reference is “load-bearing.” Classifier-independent signals may ultimately prove more reliable than any single faithfulness metric [42].